# An Imitative Reinforcement Learning Framework for Autonomous Dogfight

### v2

@@ 6. Experiments
We evaluate the proposed learning framework in the constructed simulated combat environment to validate whether the proposed framework can learn successful dogfight policies. The experiments are designed to answer the following questions:
@@
  • Can the proposed framework learn effective dogfight policy? What is the learned policy like? (Section 6.2)
@@ 6.1. Comparative Results
Table 1 and Table 2 list the parameter settings used in the experiments. Most hyper-parameters are set as the same values for all the methods, as shown in Table 1, and we tune the replay buffer size for each method to achieve their best performance. As for the neural network structure, the actor and critic networks, besides the input and output layers, consist of two hidden layers with the Relu activation function. The expert datasets for all the baseline methods are collected in AI mode with $20$ successful episodes, where the AI mode operates based on a series of predefined rules, such as directing the aircraft towards the opponent and launching missiles when a lock-on is achieved. The expert dataset contains nearly $43e$ 4 state-action pairs, provided at https://github.com/zrc0622/HIRL4UCAV. The episode terminates upon hitting the opponent aircraft down, so the lengths of successful trajectories vary.
@@
Ours(linear) $\bf{100.0\%\pm 0.0\%}$ $36.6\pm 15.8$ $\bf{100.0\%\pm 0.0\%}$ $36.4\pm 33.0$ $96.0\%\pm 2.4\%$ $-1312.5\pm 316.9$
TD3 $0.0\%\pm 0.0\%$ $-1033.3\pm 67.3$ $25.0\%\pm 43.3\%$ $-1020.9\pm 320.5$ $0.0\%\pm 0.0\%$ $-2233.3\pm 211.2$
SAC $99.0\%\pm 1.7\%$ $30.5\pm 86.3$ $98.5\%\pm 2.6\%$ $-270.6\pm 378.1$ $76.0\%\pm 16.3\%$ $-2029.7\pm 1350.1$
E-SAC $\bf{100.0\%\pm 0.0\%}$ $-248.9\pm 356.0$ $72.5\%\pm 42.0\%$ $-601.8\pm 431.8$ $73.5\%\pm 33.4\%$ $-1598.8\pm 661.1$
DSAC-v2 $0.0\%\pm 0.0\%$ $-791.1\pm 69.6$ $0.0\%\pm 0.0\%$ $-825.5\pm 34.3$ $0.0\%\pm 0.0\%$ $-1197.8\pm 137.9$
BC $\bf{100.0\%\pm 0.0\%}$ $-23.3\pm 15.4$ $73.0\%\pm 9.2\%$ $-1102.7\pm 379.9$ $65.0\%\pm 9.1\%$ $-3055.3\pm 555.9$
Expert $100.0\%\pm 0.0\%$ $9.6\pm 16.6$ $100.0\%\pm 0.0\%$ $4.3\pm 13.7$ $100.0\%\pm 0.0\%$ $-931.3\pm 154.5$
We select the best-performing results from four runs and calculate their average. The results are presented in Table 3. The data in the table is displayed in the format ”mean ± standard deviation” for clarity. The code of the proposed approach, the learned models, and recorded videos of the learned policies are available at https://github.com/zrc0622/HIRL4UCAV.
@@ 6.3. Robustness of the Learned Policy
Algorithm #$\frac{\#Hits}{\#Launches}$
@@ 7. Conclusion
This study focuses on the complex multi-dimensional decision-making problems in the UCAV dogfight. Leveraging the Harfang3D sandbox, we construct a realistic simulation platform and propose an innovative high-imitative reinforcement learning framework for dogfight policy. This method integrates expert experience with autonomous exploration mechanisms, achieving efficient and precise policy learning. Experimental results demonstrate that the constructed simulation environment serves as an effective platform for conducting reinforcement learning research in aerial combat, enabling UCAVs to handle complex, multistage dogfight tasks, including target pursuit, target lock, and missile launch in simulated air combat missions. Policies learned with the proposed framework can achieve a hit success rate of up to $100.0\%$. Compared to RL and imitation learning algorithms, the learned policies in this study demonstrate superior success rates and robustness while requiring less expert data.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grant No.62306088), and Songjiang Lab (Grant No.SL20230309).

### v3

@@ 6. Experiments
We evaluate the proposed learning framework in the constructed simulated combat environment to validate whether the proposed framework can learn successful pursuit-lock-launch policies. The experiments are designed to answer the following questions:
@@
  • Can the proposed framework learn effective pursuit-lock-launch policy? What is the learned policy like? (Section 6.2)
@@ 6.1. Comparative Results
Table 1 and Table 2 list the parameter settings used in the experiments. Most hyper-parameters are set as the same values for all the methods, as shown in Table 1, and we tune the replay buffer size for each method to achieve their best performance. As for the neural network structure, the actor and critic networks, besides the input and output layers, consist of two hidden layers with the Relu activation function. The expert datasets for all the baseline methods are collected in AI mode with $20$ successful episodes. In AI mode, the agent employs a Proportional-Integral-Derivative (PID) algorithm to regulate its attitude and velocity. The specific implementation involves the following steps: 1) Calculating the altitude deviation and heading angle deviation relative to the target; 2) Setting the autopilot parameters for target altitude and heading angle; 3) Adjusting the angles of the three wings based on the established altitude and heading parameters; 4) Iterating through these steps continuously. Through the aforementioned steps, the agent is capable of maintaining continuous orientation towards the hostile aircraft and initiating engagement upon achieving target lock. The expert dataset contains nearly $3e4$ state-action pairs, provided at https://github.com/zrc0622/HIRL4UCAV. The episode terminates upon hitting the opponent aircraft down, so the lengths of successful trajectories vary.
@@
Ours(linear) $\bf{100.0\%\pm 0.0\%}$ $36.6\pm 15.8$ $\bf{100.0\%\pm 0.0\%}$ $36.4\pm 33.0$ $96.0\%\pm 2.4\%$ $1312.5\pm 316.9$
TD3 $0.0\%\pm 0.0\%$ $1033.3\pm 67.3$ $25.0\%\pm 43.3\%$ $1020.9\pm 320.5$ $0.0\%\pm 0.0\%$ $2233.3\pm 211.2$
SAC $99.0\%\pm 1.7\%$ $30.5\pm 86.3$ $98.5\%\pm 2.6\%$ $270.6\pm 378.1$ $76.0\%\pm 16.3\%$ $2029.7\pm 1350.1$
E-SAC $\bf{100.0\%\pm 0.0\%}$ $248.9\pm 356.0$ $72.5\%\pm 42.0\%$ $601.8\pm 431.8$ $73.5\%\pm 33.4\%$ $1598.8\pm 661.1$
DSAC-v2 $0.0\%\pm 0.0\%$ $791.1\pm 69.6$ $0.0\%\pm 0.0\%$ $825.5\pm 34.3$ $0.0\%\pm 0.0\%$ $1197.8\pm 137.9$
BC $\bf{100.0\%\pm 0.0\%}$ $23.3\pm 15.4$ $73.0\%\pm 9.2\%$ $1102.7\pm 379.9$ $65.0\%\pm 9.1\%$ $3055.3\pm 555.9$
Expert $100.0\%\pm 0.0\%$ $9.6\pm 16.6$ $100.0\%\pm 0.0\%$ $4.3\pm 13.7$ $100.0\%\pm 0.0\%$ $931.3\pm 154.5$
We select the best-performing results from four runs and calculate their average. The results are presented in Table 3. The data in the table is displayed in the format ”mean ± standard deviation” for clarity. In the aforementioned experiments, we independently implemented all learning algorithms, including the algorithm framework and the GYM environment. The code of the proposed approach, the learned models, and recorded videos of the learned policies are available at https://github.com/zrc0622/HIRL4UCAV.
@@ 6.3. Robustness of the Learned Policy
Algorithm $\frac{\#Hits}{\#Launches}$
@@ 7. Conclusion
This study focuses on the complex multi-dimensional decision-making problems in the UCAV pursuit-lock-launch task. Leveraging the Harfang3D sandbox, we construct a realistic simulation platform and propose an innovative high-imitative reinforcement learning framework for pursuit-lock-launch policy. This method integrates expert experience with autonomous exploration mechanisms, achieving efficient and precise policy learning. Experimental results demonstrate that the constructed simulation environment serves as an effective platform for conducting reinforcement learning research in aerial combat, enabling UCAVs to handle complex, multistage WVR engagement tasks, including target pursuit, target lock, and missile launch in simulated air combat missions. Policies learned with the proposed framework can achieve a hit success rate of up to $100.0\%$. Compared to RL and imitation learning algorithms, the learned policies in this study demonstrate superior success rates and robustness while requiring less expert data.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grant No.62306088), the Natural Science Foundation of Heilongjiang Province (Grant No.YQ2024F007), and the Aviation Science Foundation of China (Grant No.2023Z021077001).